# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v2

@@ Abstract
Density ratio estimation is fundamental to tasks involving $f$-divergences, yet existing methods often fail under significantly different distributions or inadequately overlap supports, suffering from the density-chasm and the support-chasm problems. Additionally, prior approaches yield divergent time scores near boundaries, leading to instability. We design $\textbf{D}^{3}\textbf{RE}$, a unified framework for robust, stable and efficient density ratio estimation. We propose the dequantified diffusion bridge interpolant (DDBI), which expands support coverage and stabilizes time scores via diffusion bridges and Gaussian dequantization. Building on DDBI, the proposed dequantified Schrödinger bridge interpolant (DSBI) incorporates optimal transport to solve the Schrödinger bridge problem, enhancing accuracy and efficiency. Our method offers uniform approximation and bounded time scores in theory, and outperforms baselines empirically in mutual information and density estimation tasks. Code is available at https://github.com/Hoemr/Dequantified-Diffusion-Bridge-Density-Ratio-Estimation.git.
@@ 1 Introduction
Quantifying distributional discrepancies via $f$-divergences-defined through density ratios $r(\mathbf{x})=q_{1}(\mathbf{x})/q_{0}(\mathbf{x})$-is foundational in tasks such as domain adaptation, generative modeling, and hypothesis testing. However, directly estimating $r(\mathbf{x})$ by modeling $q_{0}$ and $q_{1}$ becomes intractable in high dimensions, motivating density ratio estimation (DRE) methods that bypass explicit density modeling (Sugiyama et al., 2012). While DRE underpins modern techniques like mutual information estimation (Colombo et al., 2021) and likelihood-free inference (Thomas et al., 2022), it struggles with a critical challenge known as the density-chasm problem, where multi-modal or divergent distributions lead to unstable ratio estimates (Rhodes et al., 2020).
@@
Existing methods like telescoping ratio estimation (TRE) (Rhodes et al., 2020) and its continuous extension DRE-$\infty$ (Choi et al., 2022) estimate density ratios via intermediate steps. TRE improves accuracy by adding more intermediate variables, but increases model complexity linearly. DRE-$\infty$ uses continuous-time score matching to avoid this, yet both face a core challenge: the support-chasm problem (see Definition 3.1), where $\mathsf{supp}(q_{0})\cap\mathsf{supp}(q_{1})$ is small or empty. This leads to inadequately overlap supports and ill-defined ratios (Srivastava et al., 2023).
@@
To address the support-chasm problem, we unify the interpolation strategies in prior works as deterministic interpolant (DI) and propose the diffusion bridge interpolant (DBI), which uses diffusion bridges to enable diverse trajectory exploration and smooth transitions between distributions. By Theorem 3.2 and Corollary 3.3, DBI expands support coverage and trajectory sets beyond existing approaches.
@@
A second challenge arises in prior methods: as $t\to 1^{-}$, the absolute time score $\mathbb{E}_{q_{t}}\left[|\partial_{t}\log q_{t}|\right]$ diverges for both DI and DBI (Theorem 3.4), leading to unstable estimations at boundary. To mitigate this, we propose Gaussian dequantization (GD), which mollifies boundary densities $q_{0}$ and $q_{1}$ via Gaussian convolution, ensuring $\mathbb{E}_{q_{t}}\left[|\partial_{t}\log q_{t}|\right]$ remains bounded over $t\in[0,1]$ (Corollary 3.8). The resulting dequantified diffusion bridge interpolant (DDBI) balances robustness and computational efficiency.
Table 1: Comparison of interpolants in this work. The second row indicates the performance benefits contributed by each module.
@@
To further reduce estimation error and improve efficiency, the dequantified Schrödinger bridge interpolant (DSBI) is derived by integrating DDBI with optimal transport rearrangement (OTR), solving the Schrödinger bridge problem (Proposition 3.6). Together, applying DDBI and DSBI to DRE leads to the Dequantified Diffusion Bridge Density Ratio Estimation ($\textbf{D}^{3}\textbf{RE}$) framework.
Experimental results show that $\text{D}^{3}\text{RE}$ improves robustness and efficiency in downstream tasks such as density ratio estimation, mutual information estimation, and density estimation. Figure 1 illustrates a comparison of interpolation strategies among DI, DBI, DDBI, and DSBI, with light blue points representing intermediate samples drawn from $q_{0}$ and $q_{1}$. Our proposed methods (DBI, DDBI, and DSBI) enable broader exploration of alternative trajectories, producing intermediate distributions with larger support compared to DI, consistent with Theorem 3.2 and Corollary 3.3.
(a) DI
@@
Figure 1: Comparison of the trajectory sets among DI, DBI, DDBI and DSBI. Our methods have larger sets of trajectories, with intermediate distributions offering larger support than those of DI.
@@
  • We propose $\text{D}^{3}\text{RE}$, the first unified framework to address both the density-chasm and support-chasm problems via uniformly approximated density ratio estimation (Proposition 3.5). Our interpolants expand both the support (Theorem 3.2) and trajectory sets (Corollary 3.3), thereby alleviating support-chasm problem.
@@
  • Experimental results demonstrate $\text{D}^{3}\text{RE}$’s superiority in density ratio estimation, mutual information estimation, and density estimation.
@@ 2 Background
Let $\mathbf{X}_{0}\sim q_{0}(\mathbf{x})$ and $\mathbf{X}_{1}\sim q_{1}(\mathbf{x})$ be random variables in $\mathbb{R}^{d}$ with set-theoretic supports $\mathsf{supp}(q_{0})$ and $\mathsf{supp}(q_{1})$. Upper cases like $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$ denote random variable. Lower cases like $\mathbf{x}_{0}$ denote the samples of random variables.
@@ 2.1 Density Ratio Estimation
Density ratio estimation (DRE) aims to estimate the density ratio $r^{\star}(\mathbf{x})=\frac{q_{1}(\mathbf{x})}{q_{0}(\mathbf{x})}$ using i.i.d. samples from both distributions. A common approach is density ratio matching via Bregman divergence minimization (Sugiyama et al., 2012), which optimizes a parameterized ratio $r_{\boldsymbol{\theta}}$ by minimizing:
$\mathrm{BD}(r_{\boldsymbol{\theta}})$ $=-\mathbb{E}_{q_{0}(\mathbf{x}_{0}),\atop q_{1}(\mathbf{x}_{1})}\left[\log\frac{1}{1+r_{\boldsymbol{\theta}}(\mathbf{x}_{0})}+\log\frac{r_{\boldsymbol{\theta}}(\mathbf{x}_{1})}{1+r_{\boldsymbol{\theta}}(\mathbf{x}_{1})}\right],$ (1)
where $\boldsymbol{\theta}$ denote the trainable parameters of $r_{\boldsymbol{\theta}}$. The minimizer of Eq. (1) satisfies $r_{\boldsymbol{\theta}^{\star}}=r^{\star}$. However, DRE suffers from the density-chasm problem (Rhodes et al., 2020).
TRE mitigates this by employing a divide-and-conquer strategy. It partitions the interval $[0,1]$ into $M\in\mathbb{Z}_{+}$ sub-intervals with endpoints $\{m/M\}_{m=0}^{M}$, constructing intermediate variables $\mathbf{X}_{m/M}\sim q_{m/M}(\mathbf{x})$ via linear interpolation:
@@
where $\eta_{m/M}$ increases from 0 to 1. The density ratio decomposes into a telescoping product:
@@
where $r_{m/M}^{\star}$ is the target intermediate density ratio, estimated by a parameterized neural network $r_{\boldsymbol{\theta}_{m}}$ with trainable parameters $\boldsymbol{\theta}_{m}$. In this case, $M$ networks must be trained. While a larger $M$ improves accuracy, it increases computational cost and may still fail to sufficiently reduce the KL divergence, $\mathrm{KL}(q_{m/M}\|q_{(m+1)/M})$, leaving the density-chasm problem unmitigated.
DRE-$\infty$ (Choi et al., 2022) extends TRE to $M\to\infty$, representing the log ratio as an integral of the time score:
@@
where $\partial_{t}\log q_{t}(\mathbf{x})\approx(\log q_{t+\Delta t}(\mathbf{x})-\log q_{t}(\mathbf{x}))/\Delta t$ with an infinitesimal gap $\Delta t=\lim_{M\to\infty}1/M$ denotes the time score. The time score model $s_{\boldsymbol{\theta}}^{t}$ approximates the time score via minimization of:
@@
where $\lambda(\cdot):[0,1]\to\mathbb{R}_{+}$ is a time-dependent weighting function, and $q(t)=\mathcal{U}[0,1]$ is the uniform distribution over the interval [0,1]. When the time score model satisfies $\partial_{t}\log q_{t}(\mathbf{x})=s_{\boldsymbol{\theta}^{\star}}^{t}(\mathbf{x},t)$, the loss function reaches its minimum value $\mathcal{L}_{1}(\boldsymbol{\theta}^{\star})$. Unlike in TRE, only one network, $s_{\boldsymbol{\theta}}^{t}$, need to be trained in DRE-$\infty$.
@@ 2.2 Denoising Diffusion Bridge Model
Denoising diffusion models (DDMs) simulate a diffusion process $\{\mathbf{X}_{t}\}_{t\in[0,1]}$, which serves as a continuous bridge between $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$. This process is described by the solution to an Itô stochastic differential equation (SDE):
@@
Conventional DDMs (Li et al., 2023; Song et al., 2020b; Li et al., 2024a; Xu et al., 2024) require either $q_{0}$ or $q_{1}$ to be a simple, tractable distribution (e.g., isotropic Gaussian), which limits their ability to bridge arbitrary complex distributions and restricts applications such as DRE.
Denoising diffusion bridge models (Zhou et al., 2023) overcome this limitation by simulating stochastic processes that interpolate between paired distributions with $\mathbf{X}_{1}$ as endpoints. These processes are derived from the SDE in Eq. (6) via Doob’s $h$-transform (Doob & Doob, 1984):
@@
where $\tilde{\mathbf{f}}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\mathbf{f}(\mathbf{X}_{t},t)+g^{2}(t)\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})$ denotes the shifted drift, $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{1}\mid\mathbf{X}_{t})$ is the $h$ function, representing the gradient of the log transition kernel from time $t$ to 1. The process explicitly depends on $\mathbf{X}_{1}$, and given $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, it forms a diffusion bridge with transition kernel $q_{t}(\mathbf{x}\mid\mathbf{X}_{0},\mathbf{X}_{1})$. A special case of the diffusion bridge, termed the Brownian bridge, arises under the conditions: $\mathbf{f}(\mathbf{X}_{t},t):=0,g(t):=1$ and $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1}):=\frac{\mathbf{X}_{1}-\mathbf{X}_{t}}{1-t}$. The Brownian bridge is defined as the solution to $\mathrm{d}\mathbf{B}_{t}=\frac{\mathbf{B}_{1}-\mathbf{B}_{t}}{1-t}\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}$, and its transition kernel is given by $q(\mathbf{B}_{t}\mid\mathbf{B}_{0},\mathbf{B}_{1})=\mathcal{N}((1-t)\mathbf{B}_{0}+t\mathbf{B}_{1},t(t-1)\mathbf{E}_{d})$.
@@ 3.1 Support-chasm Problem
Deterministic Interpolant. We summarize the interpolants used in prior works such as TRE and DRE-$\infty$ to deterministic interpolant (DI), defined as:
$\mathbf{X}_{t}=\mathbf{I}(\mathbf{X}_{0},\mathbf{X}_{1},t)=\alpha_{t}\mathbf{X}_{0}+\beta_{t}\mathbf{X}_{1},$ (8)
where $\mathbf{I}:\mathbb{R}^{d}\times\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ is an interpolant continuously differentiable in $(\mathbf{X}_{0},\mathbf{X}_{1},t)$ and the time-indexed coefficients $\alpha_{t},\beta_{t}\in C^{2}[0,1]$ are monotonic with $\alpha_{t}$ decreasing and $\beta_{t}$ increasing in $t$. They are positive definite and satisfy boundary conditions $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$, with constraints $\alpha_{t}+\beta_{t}=1$ or $\alpha_{t}^{2}+\beta_{t}^{2}=1,\forall t\in[0,1]$.
The process $\{\mathbf{X}_{t}\}_{t\in[0,1]}$, associated with Eq. (8), is termed as the deterministic interpolant (DI). Prior methods, such as TRE and DRE-$\infty$, are specific cases of DI, distinguished by their choices of $\alpha_{t}$ and $\beta_{t}$ (see Section B.1).
@@
Diffusion Bridge Interpolant. To mitigate the support-chasm problem, a Brownian bridge $\{\mathbf{B}_{t}\}_{t\in[0,1]}$ is introduced, leading to the diffusion bridge interpolant (DBI):
@@
Since $\mathbf{B}_{t}$ is a Gaussian process with zero mean and variance $t(1-t)$, the transition kernel of the DBI is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(\mathbf{I}(\mathbf{x}_{0},\mathbf{x}_{1},t),t(1-t)\gamma^{2}\mathbf{E}_{d})$, leading to an equivalent form of DBI:

### v3

@@ Abstract
Density ratio estimation is fundamental to tasks involving $f$-divergences, yet existing methods often fail under significantly different distributions or inadequately overlapping supports — the density-chasm and the support-chasm problems. Additionally, prior approaches yield divergent time scores near boundaries, leading to instability. We design $\textbf{D}^{3}\textbf{RE}$, a unified framework for robust, stable and efficient density ratio estimation. We propose the dequantified diffusion bridge interpolant (DDBI), which expands support coverage and stabilizes time scores via diffusion bridges and Gaussian dequantization. Building on DDBI, the proposed dequantified Schrödinger bridge interpolant (DSBI) incorporates optimal transport to solve the Schrödinger bridge problem, enhancing accuracy and efficiency. Our method offers uniform approximation and bounded time scores in theory, and outperforms baselines empirically in mutual information and density estimation tasks. Code is available at https://github.com/Hoemr/Dequantified-Diffusion-Bridge-Density-Ratio-Estimation.git.
@@ 1 Introduction
Quantifying distributional discrepancies via $f$-divergences-defined through density ratios $r(\mathbf{x})=q_{1}(\mathbf{x})/q_{0}(\mathbf{x})$ is foundational in tasks such as domain adaptation, generative modeling, and hypothesis testing. However, directly estimating $r(\mathbf{x})$ by modeling $q_{0}$ and $q_{1}$ becomes intractable in high dimensions, motivating density ratio estimation (DRE) methods that bypass explicit density modeling (Sugiyama et al., 2012). While DRE underpins modern techniques like mutual information estimation (Colombo et al., 2021) and likelihood-free inference (Thomas et al., 2022), it struggles with a critical challenge known as the density-chasm problem, where multi-modal or divergent distributions lead to unstable ratio estimates (Rhodes et al., 2020).
@@
Existing methods like telescoping ratio estimation (TRE) (Rhodes et al., 2020) and its continuous extension DRE-$\infty$ (Choi et al., 2022) estimate density ratios via intermediate steps. TRE improves accuracy by adding more intermediate variables, but increases model complexity linearly. DRE-$\infty$ uses continuous-time score matching to avoid this, yet both face a core challenge in the support-chasm problem (see Definition 3.1), where $\mathsf{supp}(q_{0})\cap\mathsf{supp}(q_{1})$ is small or empty. This leads to inadequately overlapping supports and ill-defined ratios (Srivastava et al., 2023).
@@
To address the support-chasm problem, we unify the interpolation strategies in prior works as deterministic interpolants (DI) and propose the diffusion bridge interpolant (DBI), which uses diffusion bridges to enable diverse trajectory exploration and smooth transitions between distributions. By Theorem 3.2 and Corollary 3.3, DBI expands support coverage and trajectory sets beyond existing approaches.
@@
A second challenge arises in prior methods: As $t\to 1^{-}$, the absolute time score $\mathbb{E}_{q_{t}}\left[|\partial_{t}\log q_{t}|\right]$ diverges for both DI and DBI (Theorem 3.4), leading to unstable estimations at the boundary. To mitigate this, we propose Gaussian dequantization (GD), which addresses boundary densities $q_{0}$ and $q_{1}$ via Gaussian convolution, ensuring $\mathbb{E}_{q_{t}}\left[|\partial_{t}\log q_{t}|\right]$ remains bounded over $t\in[0,1]$ (Corollary 3.8). The resulting dequantified diffusion bridge interpolant (DDBI) balances robustness and computational efficiency.
To further reduce estimation error and improve efficiency, the dequantified Schrödinger bridge interpolant (DSBI) is derived by integrating DDBI with optimal transport rearrangement (OTR), solving the Schrödinger bridge problem (Proposition 3.6). Together, applying DDBI and DSBI to DRE leads to the Dequantified Diffusion Bridge Density Ratio Estimation ($\textbf{D}^{3}\textbf{RE}$) framework. We summarize these developments in Table 1.
Experimental results show that $\text{D}^{3}\text{RE}$ improves robustness and efficiency in downstream tasks such as density ratio estimation, mutual information estimation, and density estimation. Figure 1 illustrates a comparison of interpolation strategies among DI, DBI, DDBI, and DSBI, with light blue points representing intermediate samples drawn from $q_{0}$ and $q_{1}$. Our proposed methods (DBI, DDBI, and DSBI) enable broader exploration of alternative trajectories, producing intermediate distributions with larger support compared to DI, consistent with Theorem 3.2 and Corollary 3.3 below.
Table 1: Comparison of advantages of interpolants in this work.
@@
(a) DI (previous)
@@
Figure 1: Trajectory sets comparison among DI, DBI, DDBI and DSBI. Our methods yield broader trajectory sets, with intermediate distributions exhibiting wider support than those of DI. The entropically regularized transport losses for subfigures (a-d) are $44.17$, $31.14$, $31.15$, and $8.26$, respectively (See Equation 17 for details.). Lower loss indicates increased path diversity.
@@
  • We propose $\text{D}^{3}\text{RE}$, the first unified framework to address both the density-chasm and support-chasm problems via uniformly approximated density ratio estimation (Proposition 3.5). Our interpolants expand both the support (Theorem 3.2) and trajectory sets (Corollary 3.3), alleviating the support-chasm problem.
@@
  • Experiments demonstrate $\text{D}^{3}\text{RE}$’s superiority in density ratio estimation, mutual information estimation, and density estimation.
@@ 2 Background
Let $\mathbf{X}_{0}\sim q_{0}(\mathbf{x})$ and $\mathbf{X}_{1}\sim q_{1}(\mathbf{x})$ be random variables in $\mathbb{R}^{d}$ with set-theoretic supports $\mathsf{supp}(q_{0})$ and $\mathsf{supp}(q_{1})$. Upper cases $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$ denote random variable, while lower cases $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ denote the samples of random variables.
@@ 2.1 Density Ratio Estimation
Density ratio estimation (DRE) aims to estimate the density ratio $r^{\star}(\mathbf{x})=\frac{q_{1}(\mathbf{x})}{q_{0}(\mathbf{x})}$ using i.i.d. samples from both distributions. A common approach is density ratio matching via Bregman divergence minimization (Sugiyama et al., 2012), which optimizes a parameterized ratio $r_{\boldsymbol{\theta}}$ by minimizing
$\mathrm{BD}(r_{\boldsymbol{\theta}})=-\mathbb{E}_{\begin{subarray}{c}q_{0}(\mathbf{x}_{0})\\ q_{1}(\mathbf{x}_{1})\end{subarray}}\left[\log\frac{1}{1+r_{\boldsymbol{\theta}}(\mathbf{x}_{0})}+\log\frac{r_{\boldsymbol{\theta}}(\mathbf{x}_{1})}{1+r_{\boldsymbol{\theta}}(\mathbf{x}_{1})}\right]$,
where $\boldsymbol{\theta}$ denotes the trainable parameters of $r_{\boldsymbol{\theta}}$. The minimizer of Eq. (1) satisfies $r_{\boldsymbol{\theta}^{\star}}=r^{\star}$. However, DRE suffers from the density-chasm problem (Rhodes et al., 2020).
TRE mitigates this by employing a divide-and-conquer strategy. It partitions the interval $[0,1]$ into $M\in\mathbb{Z}_{+}$ sub-intervals with endpoints $\{m/M\}_{m=0}^{M}$, constructing intermediate variables $\mathbf{X}_{m/M}\sim q_{m/M}(\mathbf{x})$ via linear interpolation
@@
where $\eta_{m/M}$ increases from 0 to 1. The density ratio decomposes into a telescoping product
@@
where $r_{m/M}^{\star}$ is the target intermediate density ratio, estimated by a parameterized neural network $r_{\boldsymbol{\theta}_{m}}$ with trainable parameters $\boldsymbol{\theta}_{m}$. In this case, $M$ networks must be trained. While a larger $M$ improves accuracy, it increases computational cost and may still fail to sufficiently reduce the KL divergence, $\mathrm{KL}(q_{m/M}\|q_{(m+1)/M})$, leaving the density-chasm problem unaddressed.
DRE-$\infty$ (Choi et al., 2022) extends TRE to $M\to\infty$, representing the log ratio as an integral of the time score
@@
where $\partial_{t}\log q_{t}(\mathbf{x})\approx(\log q_{t+\Delta t}(\mathbf{x})-\log q_{t}(\mathbf{x}))/\Delta t$ with an infinitesimal gap $\Delta t=\lim_{M\to\infty}1/M$ denotes the time score. The time score model $s_{\boldsymbol{\theta}}^{t}$ approximates the time score via minimization of
@@
where $\lambda(\cdot):[0,1]\to\mathbb{R}_{+}$ is a time-dependent weighting function, and $q(t)=\mathcal{U}[0,1]$ is the uniform distribution over the interval [0,1]. When the time score model satisfies $\partial_{t}\log q_{t}(\mathbf{x})=s_{\boldsymbol{\theta}^{\star}}^{t}(\mathbf{x},t)$, the loss function reaches its minimum value $\mathcal{L}_{1}(\boldsymbol{\theta}^{\star})$. Unlike in TRE, only one network, $s_{\boldsymbol{\theta}}^{t}$, needs to be trained in DRE-$\infty$.
@@ 2.2 Denoising Diffusion Bridge Model
Denoising diffusion models (DDMs) simulate a diffusion process $\{\mathbf{X}_{t}\}_{t\in[0,1]}$, which serves as a continuous bridge between $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$. This process is described by the solution to an Itô stochastic differential equation (SDE)
@@
Conventional DDMs (Li et al., 2023; Song et al., 2021; Li et al., 2024a; Xu et al., 2024) require either $q_{0}$ or $q_{1}$ to be a simple, tractable distribution (e.g., isotropic Gaussian), which limits their ability to bridge arbitrary complex distributions and restricts applications such as DRE.
Denoising diffusion bridge models (Zhou et al., 2024a) overcome this limitation by simulating stochastic processes that interpolate between paired distributions with $\mathbf{X}_{1}$ as endpoints. These processes are derived from the SDE in Eq. (6) via Doob’s $h$-transform (Doob & Doob, 1984)
@@
where $\tilde{\mathbf{f}}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\mathbf{f}(\mathbf{X}_{t},t)+g^{2}(t)\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})$ denotes the shifted drift and $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{1}\mid\mathbf{X}_{t})$ is the $h$ function representing the gradient of the log transition kernel from time $t$ to 1. The process explicitly depends on $\mathbf{X}_{1}$ and, given $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, it forms a diffusion bridge with transition kernel $q_{t}(\mathbf{x}\mid\mathbf{X}_{0},\mathbf{X}_{1})$. A special case of the diffusion bridge, termed the Brownian bridge, arises under the conditions $\mathbf{f}(\mathbf{X}_{t},t):=0,g(t):=1$ and $\mathbf{h}(\mathbf{X}_{t},t,\mathbf{X}_{1}):=\frac{\mathbf{X}_{1}-\mathbf{X}_{t}}{1-t}$. The Brownian bridge is defined as the solution to $\mathrm{d}\mathbf{B}_{t}=\frac{\mathbf{B}_{1}-\mathbf{B}_{t}}{1-t}\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}$ and its transition kernel is given by $q(\mathbf{B}_{t}\mid\mathbf{B}_{0},\mathbf{B}_{1})=\mathcal{N}((1-t)\mathbf{B}_{0}+t\mathbf{B}_{1},t(t-1)\mathbf{E}_{d})$.
@@ 3.1 Support-chasm Problem
Deterministic Interpolant. We summarize the interpolants used in prior works, such as TRE and DRE-$\infty$, to deterministic interpolant (DI), defined as
$\mathbf{I}(\mathbf{X}_{0},\mathbf{X}_{1},t)=\alpha_{t}\mathbf{X}_{0}+\beta_{t}\mathbf{X}_{1},$ (8)
where $\mathbf{I}:\mathbb{R}^{d}\times\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ is an interpolant continuously differentiable in $(\mathbf{X}_{0},\mathbf{X}_{1},t)$ and the time-indexed coefficients $\alpha_{t},\beta_{t}\in C^{2}[0,1]$ are monotonic with $\alpha_{t}$ decreasing and $\beta_{t}$ increasing in $t$. They are strictly positive and satisfy boundary conditions $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$, with constraints $\alpha_{t}+\beta_{t}=1$ or $\alpha_{t}^{2}+\beta_{t}^{2}=1,\forall t\in[0,1]$.
Prior methods, such as TRE (Rhodes et al., 2020) and DRE-$\infty$ (Choi et al., 2022), are specific cases of DI distinguished by their choices of $\alpha_{t}$ and $\beta_{t}$ (see Section B.1).
@@
Diffusion Bridge Interpolant. To mitigate the support-chasm, we introduce a Brownian bridge $\{\mathbf{B}_{t}\}_{t\in[0,1]}$, leading to the diffusion bridge interpolant (DBI)
@@
Since $\mathbf{B}_{t}$ is a Gaussian process with zero mean and variance $t(1-t)$, the transition kernel of the DBI is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(\mathbf{I}(\mathbf{x}_{0},\mathbf{x}_{1},t),t(1-t)\gamma^{2}\mathbf{E}_{d})$, leading to an equivalent form of DBI